OpenStudy (anonymous):

Four 2 kg masses are located at the corners of a square of sides 25 cm. What is the moment of inertia of the array about an axis through one of the masses and perpendicular to the plane of the four masses?

OpenStudy (anonymous):

 moment of inertia I = mr^2 for this system, I = m1 r1^2 + m2 r2^2 + m3 r3^2 + m4 r4^2 since m1 = m2 = m3 = m4 = 2 I = 2 ( r1^2 + r2^2 + r3^2+ r4^2 ) I = 2 ( r1^2 + r2^2 + r3^2 + 0 )...........since m4 lies on the axis itself, its distance from the axis is zero. now, find r1, r2, r3 and u'll have ur answer.